OpenStudy (nathanjhw):

The position of a particle moving in a straight line is given by s(t)=e^12t cos (5t) for t > 0 where t is in seconds. If the particle changes direction at time T seconds, then T must satisfy the equation:

OpenStudy (nathanjhw):

tan(5T ) = 2.4 tan(5T ) = 4.8 cos(5T ) = 0

OpenStudy (nathanjhw):

@dan815

OpenStudy (nathanjhw):

Yes

OpenStudy (nathanjhw):

I'm still confused though.

OpenStudy (nathanjhw):

@dan815

OpenStudy (anonymous):

Changing direction would imply that the derivative is 0 or undefined. Basically T has to be a critical point.

OpenStudy (nathanjhw):

So then the answer would be cos(5T)=0?

OpenStudy (nathanjhw):

@wio

OpenStudy (anonymous):

\[ s'(t) = 12e^{12t}\cos(5t) - 5e^{12t}\sin(5t) = e^{12t}\left(\frac{12}{5} -\tan(5t)\right) \]If we want \(s'(t)=0\), considering that \(e^{12t}>0\) for all \(t\), then we want \[ \frac{12}{5} -\tan(5t)=0 \implies \tan(5t) = \frac{12}{5} = 2.4 \]

OpenStudy (nathanjhw):

Thank you very much @wio !

OpenStudy (anonymous):

Not sure if you learned anything

OpenStudy (anonymous):

I made a mistake anyway

OpenStudy (anonymous):

\[ s'(t) = 12e^{12t}\cos(5t) - 5e^{12t}\sin(5t) = 5e^{12t}\cos(5t)\left(\frac{12}{5} -\tan(5t)\right) \]

OpenStudy (anonymous):

So there are two potential equations to satisfy.

OpenStudy (dan815):

Oops ignore what i write >_> i am tired.. i thought changing direction meant going all the way back